Please help me!

Question:

Brenda baked a delicious batch of brownies in a pan.

1. She divided up the brownies into 15 equal sized rectangles. Draw the rectangles to show how the
brownie would be cut in the pan.

2. The pan has dimensions 12 inches by 8 inches. What are the dimensions of one of your brownies?
______________ inches by ______________ inches

3. She invites three friends over to share the brownies among the four of them. They all ate the same amount and all the brownies were eaten. How much did each person eat?
For a pan of Brenda’s brownies, she needs 3/8 of a cup of flour and 1/3 of a cup of cocoa powder.

4. Brenda has a package of flour that contains 2 ¾ cups of flour. How many pans of brownies can she make?

5. If she also has 2 ½ cups of cocoa powder in a package, which ingredient will she run out of first? Explain how you know for sure.

Answer by jim_thompson5910(35256)

Brenda baked a delicious batch of brownies in a pan.

1. She divided up the brownies into 15 equal sized rectangles. Draw the rectangles to show how the
brownie would be cut in the pan.


Divide the pan into a 3x5 batch of rectangles. So you'll have 3 rows and 5 columns, or vice versa.

-----------------------------------------------------------------------------------------------------------------
2. The pan has dimensions 12 inches by 8 inches. What are the dimensions of one of your brownies?
______________ inches by ______________ inches


If you have 3 brownies run across the 12 inch dimension, then (12 inches)/(3 brownies) = 4 inches is one dimension of each brownie
If you have 5 brownies run across the 8 inch dimension, then (8 inches)/(5 brownies) = 1.6 inches is the other dimension of each brownie

So each brownie is 4 inches by 1.6 inches

Note: 1.6 = 1 & 3/5

-----------------------------------------------------------------------------------------------------------------

3. She invites three friends over to share the brownies among the four of them. They all ate the same amount and all the brownies were eaten. How much did each person eat?


There are 15 brownies to divide among 4 people

15/4 = 4.75

So each person ate 4.75 brownies

Note: 4 & 3/4 = 4.75

------------------------------------------------------------------------------------------------------------------
For a pan of Brenda’s brownies, she needs 3/8 of a cup of flour and 1/3 of a cup of cocoa powder.
4. Brenda has a package of flour that contains 2 ¾ cups of flour. How many pans of brownies can she make?


x = number of pans

(3/8)*x = 2 & 3/4
(3/8)*x = 2 + 3/4
(3/8)*x = 2*(4/4) + 3/4
(3/8)*x = 8/4 + 3/4
(3/8)*x = 11/4
(8/3)*(3/8)*x = (8/3)*(11/4)
x = 88/12
x = 22/3
x = (21+1)/3
x = 21/3+1/3
x = 7+1/3
x = 7 & 1/3

She can make 7 & 1/3 pans of brownies. Round down to get 7 whole pans. She will have leftover flour, but not enough to make another whole pan (to bring it up to 8)

Answer: 7 whole pans.

------------------------------------------------------------------------------------------------------------------
5. If she also has 2 ½ cups of cocoa powder in a package, which ingredient will she run out of first? Explain how you know for sure.


total amount of cocoa powder available = (amount of cocoa powder per pan)*(number of pans)
2+1/2 = (1/3)*(number of pans)
4/2+1/2 = (1/3)*(number of pans)
5/2 = (1/3)*(number of pans)
(5/2)*3 = number of pans
15/2 = number of pans
number of pans = 7+1/2

We can make 7&1/3 pans of brownies with the amount of flour given.
We can make 7&1/2 pans of brownies with the amount of cocoa powder given.

Therefore, we will run out of flour first (because 7&1/3 is less than 7&1/2). But it doesn't matter because 7 is smaller than both of these, so we can easily make 7 pans with leftover ingredients of both types.